# Non-abelian Hodge correspondence over singular Kähler spaces

### v1

@@ 3.1.2. Bando-Siu’s methods
as $\epsilon\rightarrow 0$ in $C^{\infty}_{loc}(X\setminus\Sigma)$, where $\Omega_{\epsilon}=\omega_{1,\epsilon}\wedge\cdots\wedge\omega_{n-2,\epsilon}$ and $\Omega=\omega_{1}\wedge\cdots\wedge\omega_{n-2}$. Combining (3.3), Remark 2.14, Lemma 3.15 and the Fatou Lemma, we have
@@
$=$ $-8\pi^{2}\lim\limits_{\epsilon\rightarrow 0}\mathrm{ch}_{2}(E_{\mathrm{orb}})\cdot[\omega_{1,\epsilon}]\cdots[\omega_{n-2,\epsilon}]=\lim\limits_{\epsilon\rightarrow 0}\int_{Y_{\mathrm{orb}}}\operatorname{tr}(F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}\wedge F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}})\wedge\Omega_{\epsilon}$
$=$ $\lim\limits_{\epsilon\rightarrow 0}\int_{Y_{\mathrm{orb}}}\operatorname{tr}\left((F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}-\frac{\Lambda_{\omega_{\epsilon}}F_{H_{\mathrm{orb},\epsilon}(t)}}{n}\cdot\gamma_{\epsilon}\omega_{0,\epsilon})\wedge(F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}-\frac{\Lambda_{\omega_{\epsilon}}F_{H_{\mathrm{orb},\epsilon}(t)}}{n}\cdot\gamma_{\epsilon}\omega_{0,\epsilon})\right)\wedge\Omega_{\epsilon}$
$-\frac{1}{n^{2}}\int_{Y_{\mathrm{orb}}}|\Lambda_{\omega_{\epsilon}}F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}|^{2}\gamma_{\epsilon}^{2}\omega_{0,\epsilon}^{2}\wedge\Omega_{\epsilon}$
@@
Similarly, let $t\rightarrow+\infty$, it follows from Theorem 3.5, (3.12) and (3.11) that
@@
Since $F_{H_{\infty},\theta_{X_{\mathrm{reg}}}}\wedge\omega_{1}\wedge\cdots\wedge\omega_{n-2}=\Lambda_{\omega}F_{H_{\infty},\theta_{X_{\mathrm{reg}}}}\cdot\omega^{n-1}=0$ outside $\Sigma$, it follows from Lemma 3.7 again that $F_{H_{\infty},\theta_{X_{\mathrm{reg}}}}\equiv 0$ outside $\Sigma$.
To complete the proof, it remains to prove Bogomolov-Gieseker inequality (3.1). Recall [110, Corollary 1.3] that for a $(\omega_{0},\cdots,\omega_{n-2})$-semistable reflexive Higgs sheaf $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{\mathrm{reg}})$,
@@
with the induced Higgs sheaf $\theta_{\mathcal{G}_{X_{\mathrm{reg}}}}$ is $(\omega_{0},\cdots,\omega_{n-2})$-semistable. It follows from [110, Lemma 2.45] that $\widehat{c}_{1}(\mathcal{G}_{X})=0$ and
(3.13) $\widehat{\Delta}(\mathcal{E}_{X})=\frac{1}{\operatorname{rank}(\mathcal{E}_{X})^{2}}\widehat{\Delta}(\mathcal{G}_{X})=-\frac{1}{\operatorname{rank}(\mathcal{E}_{X})}\widehat{\mathrm{ch}}_{2}(\mathcal{G}_{X})$
Then the Bogomolov-Gieseker inequality (3.1) can be concluded by applying the above argument to $(\mathcal{G}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$.
3.2. $C_{loc}^{\infty}$-convergence of $H_{\mathrm{orb},\epsilon}(t)\rightarrow H(t)$
Recall that $H_{\mathrm{orb},\epsilon}(t)$ is the long time solution of the Hermitian-Yang-Mills flow (3.8) on the Higgs orbi-bundle $(E_{\mathrm{orb}},\theta_{\mathrm{orb}})$ with the fixed smooth initial metric $\hat{H}_{\mathrm{orb}}$ and with respect to the Hermitian metric $\omega_{\mathrm{orb},\epsilon}$. We shall prove the following statement.
@@ 3.2. $C_{loc}^{\infty}$-convergence of $H_{\mathrm{orb},\epsilon}(t)\rightarrow H(t)$
By choosing subsequences, $H_{\mathrm{orb},\epsilon}(t)$ converges to $\{f_{\mu}^{*}H(t)\}$ in $C_{loc}^{\infty}(Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(\Sigma))$, where $H(t)$ is a smooth solution of the evolution equation (3.9) on $(X\setminus\Sigma)\times[0,+\infty)$ in $C_{loc}^{\infty}$-topology as $\epsilon\rightarrow 0$. Furthermore, $H(t)$ satisfies:
(3.14) $\int_{X\setminus\Sigma}|\Phi(H(t),\omega)|_{H(t)}\frac{\omega^{n}}{n!}\leq\int_{X\setminus\Sigma}|\Phi(\hat{H},\omega)|_{\hat{H}}\frac{\omega^{n}}{n!}\leq\hat{C}_{1},$
@@
(3.15) $|\Phi(H(t+s),\omega)|_{H(t+t_{0})}(x)\leq\tilde{C}_{1}(\frac{1}{s^{2n+1}}+1)\int_{X\setminus\Sigma}|\Phi(H(t),\omega)|_{H(t)}\frac{\omega^{n}}{n!}$
for all $x\in X\setminus\Sigma$, $s>0$ and $t\geq 0$, where $\Phi(H,\omega)=\sqrt{-1}\Lambda_{\omega}(F_{H}+[\theta,\theta^{\ast H}])-\lambda_{\omega}\operatorname{Id}_{E}.$
3.2.1. Uniform Sobolev inequality for $\omega_{i,\epsilon}$
@@
(3.16) $(\int_{Y_{\mathrm{orb}}}|u|^{2q}\frac{\omega_{\mathrm{orb},k,\epsilon}^{n}}{n!})^{\frac{1}{q}}\leq C_{S}\left(\int_{Y_{\mathrm{orb}}}(|du|_{\omega_{\mathrm{orb},k,\epsilon}}^{2}+|u|^{2})\frac{\omega_{\mathrm{orb},k,\epsilon}^{n}}{n!}\right)$
@@
First, let us recall the uniform Sobolev inequality proved in [45, Section 4] (see also [91, Section 2.3] for general compact Hermitian manifolds).
Lemma 3.10.
@@
$e^{F_{\omega}}\geq\gamma,\ \|e^{F_{\omega}}\|_{L^{1+\delta}}\leq K,\ [\omega]\cdot[\theta_{\widehat{Y}}]\leq A\text{ and }[\omega]^{n}\geq A^{-1}$
where $F_{\omega}:=\log(\frac{1}{V_{\omega}}\frac{\omega^{n}}{\theta_{\widehat{Y}}^{n}})$ for $V_{\omega}:=[\omega]^{n}$. For any $q\in(1,\frac{n}{n-1})$, there exists some constant $C=C(\gamma,A,K,\delta,q)>0$ such that for any $\omega\in\mathcal{K}(\widehat{Y},\theta_{\widehat{Y}},\gamma,A,K,\delta)$, we have that
$(\int_{\widehat{Y}}|u|^{2q}\omega^{n})^{\frac{1}{q}}\leq C_{S}\left(\int_{\widehat{Y}}(|du|_{\omega}^{2}+|u|^{2})\frac{\omega^{n}}{n!}\right)$
@@
Building on Lemma 3.10, Proposition 3.9 can be directly concluded by the argument using smooth approximation from [28, Section 3]. For reader’s convenience, we include a brief explanation.
@@
For simplicity, we omit the subscript ‘k’ at the moment. Suppose that $\pi_{\mu}:U_{\mu}\rightarrow U_{\mu}/G_{\mu}$ is étale outside a proper analytic subspace $Z$. Recall Proposition 2.19 that there exists some smooth Kähler form $\omega_{0}$ on the underlying space $Y$ such that $\omega_{Y_{\mathrm{orb}}}:=\{\pi_{\mu}^{*}\omega_{0}+\sqrt{-1}\partial\bar{\partial}\varphi_{\mu}\}$ for some $G_{\mu}$-invariant smooth function $\varphi_{\mu}$ on every chart $U_{\mu}$. Then $\varphi_{\mu}$ induces a bounded continous function $\varphi$ on $Y$ that is smooth outside $Z$ and in particular, $\omega_{Y_{\mathrm{orb}}}=\{\pi_{\mu}^{*}\omega_{Y}\}$ for $\omega_{Y}=\omega_{0}+\sqrt{-1}\partial\bar{\partial}\varphi$. Thus, $\omega_{\mathrm{orb},\epsilon}:=\{f_{\mathrm{orb}}^{*}\omega+\epsilon(\pi_{\mu}\omega_{0}+\sqrt{-1}\partial\bar{\partial}\varphi_{\mu})\}$ can also be viewed as a Kähler current $\omega_{\epsilon}=\omega+\epsilon\omega_{0}+\epsilon\sqrt{-1}\partial\bar{\partial}\varphi$ on the underlying complex space $Y$ that is smooth outside $Z$. Let $f:\widehat{Y}\rightarrow Y$ be a resolution of singularities of $Y$ and $D$ be the exceptional divisor of $f\circ g$, we use the notation of Proposition 3.9 and Lemma 3.10.
As in [45, Section 7], the basic strategy is considering an approximation of $f^{*}\omega_{\epsilon}$ by a family $\{\omega_{\epsilon,j}\}_{j}$ of smooth Kähler metrics on $\widehat{Y}$. However, the situation here is slightly different to [45, Section 7] because $\log(\frac{f^{*}\omega_{\epsilon}^{n}}{\theta_{Y}^{n}})$ does not have log type analytic subvarieties (see [45, Definition 7.2]) in general, and is indeed the sum of a function $F_{\log}$ with log type singularities and a continuous function $G_{\epsilon}$ which is smooth outside $D$ (see [28, Lemma 3.9]). The main consideration in [28, Section 3] is addressing this issue. According to [28, Lemma 3.12, Lemma 3.13 and Lemma 3.15], there exists a sequence $\omega_{\epsilon,j}$ of Kähler metrics on $Y$ such that $\omega_{j,\epsilon}\in\mathcal{K}(Y,\theta_{Y},\gamma,A,K,\delta)$ for some $A,K,\delta$ and $\gamma$ independent of $j,\epsilon$, and for any integer $k\geq 0$, we have $\omega_{\epsilon,j}\rightarrow f^{*}\omega_{\epsilon}$ in $C_{loc}^{k}(Y\setminus D)$. Then by applying the same argument of [45, Pages 34] and Lemma 3.10, the uniform Sobolev inequality (3.16) for $\omega_{\epsilon}$ can be proved because $\mu_{\alpha}$ is unramified outside $Z$. ∎
It’s easy to see that there exists some constant $C_{k}>0$ independent of $\epsilon$ such that
@@
from the construction (see (3.5) and (3.6)). We have following uniform Sobolev inequalities for the Hermitian metrics $\omega_{\mathrm{orb},\epsilon}$ by applying Proposition 3.9 to $q=\frac{2n}{2n-1}$.
@@
(3.18) $(\int_{Y_{\mathrm{orb}}}|u|^{\frac{4n}{2n-1}}\frac{\omega_{\mathrm{orb},\epsilon}^{n}}{n!})^{\frac{2n-1}{2n}}\leq C_{S}\Big(\int_{Y_{\mathrm{orb}}}(|du|^{2}_{\omega_{\mathrm{orb},\epsilon}}+|u|^{2})\frac{\omega_{\mathrm{orb},\epsilon}^{n}}{n!}\Big)$

### v2

@@ 3.1.2. Bando-Siu’s methods
as $\epsilon\rightarrow 0$ in $C^{\infty}_{loc}(X\setminus\Sigma)$, where $\Omega_{\mathrm{orb},\epsilon}=\omega_{\mathrm{orb},1,\epsilon}\wedge\cdots\wedge\omega_{\mathrm{orb},n-2,\epsilon},\Omega=\omega_{1}\wedge\cdots\wedge\omega_{n-2}$ and $\gamma_{0}=\frac{\omega^{n}}{\omega_{0}^{2}\wedge\omega_{1}\cdots\wedge\omega_{n-2}}$. Combining (3.3) and Remark 2.14, and applying the Fatou Lemma, we have
@@
$=$ $-8\pi^{2}\lim\limits_{\epsilon\rightarrow 0}\mathrm{ch}_{2}^{\mathrm{orb}}(E_{\mathrm{orb}})\cdot[\omega_{\mathrm{orb},1,\epsilon}]\cdots[\omega_{\mathrm{orb},n-2,\epsilon}]=\lim\limits_{\epsilon\rightarrow 0}\int_{Y_{\mathrm{orb}}}\operatorname{tr}(F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}\wedge F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}})\wedge\Omega_{\mathrm{orb},\epsilon}$
$=$ $\lim\limits_{\epsilon\rightarrow 0}\int_{Y_{\mathrm{orb}}}\bigg(\operatorname{tr}\big((F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}-\frac{\Lambda_{\omega_{\mathrm{orb},\epsilon}}F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}}{n}\cdot\gamma_{\mathrm{orb},\epsilon}\omega_{\mathrm{orb},0,\epsilon})\wedge(F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -\frac{\Lambda_{\omega_{\mathrm{orb},\epsilon}}F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}}{n}\cdot\gamma_{\mathrm{orb},\epsilon}\omega_{\mathrm{orb},0,\epsilon})\big)-\frac{1}{n^{2}}|\Lambda_{\omega_{\mathrm{orb},\epsilon}}F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}|^{2}\gamma_{\mathrm{orb},\epsilon}^{2}\omega_{\mathrm{orb},0,\epsilon}^{2}\bigg)\wedge\Omega_{\mathrm{orb},\epsilon}$
@@
Similarly, as $t\rightarrow+\infty$, it follows from Theorem 3.5 and the equation (3.12) that
@@
Since (3.12) implies $F_{H_{\infty},\theta_{X_{\mathrm{reg}}}}\wedge\omega_{0}\wedge\cdots\wedge\omega_{n-2}=0$ outside $\Sigma$, Lemma 3.7 then yields $F_{H_{\infty},\theta_{X_{\mathrm{reg}}}}\equiv 0$ outside $\Sigma$.
To complete the proof, it remains to prove Bogomolov-Gieseker inequality (3.1). Recall from [113, Corollary 1.3] that for an $(\omega_{0},\cdots,\omega_{n-2})$-semistable reflexive Higgs sheaf $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{\mathrm{reg}})$,
@@
with the induced Higgs sheaf $\theta_{\mathcal{G}_{X_{\mathrm{reg}}}}$ is $(\omega_{0},\cdots,\omega_{n-2})$-semistable. It follows from [113, Lemma 2.45] that $\widehat{c}_{1}(\mathcal{G}_{X})=0$ and
(3.13) $\widehat{\Delta}(\mathcal{E}_{X})=\frac{1}{\operatorname{rank}(\mathcal{E}_{X})^{2}}\widehat{\Delta}(\mathcal{G}_{X})=-\frac{1}{\operatorname{rank}(\mathcal{E}_{X})}\widehat{\mathrm{ch}}_{2}(\mathcal{G}_{X}).$
Then the Bogomolov-Gieseker inequality (3.1) can be concluded by applying the proceeding argument to $(\mathcal{G}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$.
3.2. $C_{loc}^{\infty}$-convergence of $H_{\mathrm{orb},\epsilon}(t)\rightarrow H_{\mathrm{orb}}(t)$
We adopt the notation of Section 3.1 throughout Sections 3.2 and 3.3. Recall that $H_{\mathrm{orb},\epsilon}(t)$ is the long time solution of the Hermitian-Yang-Mills flow (3.8) on the Higgs orbi-bundle $(E_{\mathrm{orb}},\theta_{\mathrm{orb}})$ with the fixed smooth initial metric $\hat{H}_{\mathrm{orb}}$ and with respect to the Hermitian metric $\omega_{\mathrm{orb},\epsilon}$. We shall prove the following result.
@@ 3.2. $C_{loc}^{\infty}$-convergence of $H_{\mathrm{orb},\epsilon}(t)\rightarrow H_{\mathrm{orb}}(t)$
By choosing subsequences, $H_{\mathrm{orb},\epsilon}(t)$ converges to $H_{\mathrm{orb}}(t)=\{f_{\alpha}^{*}H(t)\}$ in $C_{loc}^{\infty}(Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(\Sigma))$ as $\epsilon\rightarrow 0$, where $H(t)$ is a smooth solution of the evolution equation (3.9) on $(X\setminus\Sigma)\times[0,+\infty)$. Furthermore, $H(t)$ satisfies:
(3.14) $\int_{X\setminus\Sigma}|\Phi(H(t),\omega)|_{H(t)}\frac{\omega^{n}}{n!}\leq\int_{X\setminus\Sigma}|\Phi(\hat{H},\omega)|_{\hat{H}}\frac{\omega^{n}}{n!}\leq\widetilde{C}_{1},$
@@
(3.15) $|\Phi(H(t+s),\omega)|_{H(t+s)}(x)\leq\tilde{C}_{2}(\frac{1}{s^{2n+1}}+1)\int_{X\setminus\Sigma}|\Phi(H(t),\omega)|_{H(t)}\frac{\omega^{n}}{n!}$
for all $x\in X\setminus\Sigma$, $s>0$ and $t\geq 0$, where $\Phi(H,\omega)=\sqrt{-1}\Lambda_{\omega}(F_{H}+[\theta,\theta^{\ast H}])-\lambda_{\omega}\operatorname{Id}_{E}$, $\widetilde{C}_{1}$ and $\tilde{C}_{2}$ are positive constants.
3.2.1. Uniform Sobolev inequality for $\omega_{\mathrm{orb},k,\epsilon}$
@@
(3.16) $(\int_{Y_{\mathrm{orb}}}|u|^{2q}\frac{\omega_{\mathrm{orb},k,\epsilon}^{n}}{n!})^{\frac{1}{q}}\leq C_{S}\int_{Y_{\mathrm{orb}}}(|du|_{\omega_{\mathrm{orb},k,\epsilon}}^{2}+|u|^{2})\frac{\omega_{\mathrm{orb},k,\epsilon}^{n}}{n!}$
@@
Recall the uniform Sobolev inequality proved in [45, Section 4] (see also [93, Section 2.3] for general compact Hermitian manifolds).
Theorem 3.10.
@@
$e^{F_{\omega}}\geq\gamma,\quad\|e^{F_{\omega}}\|_{L^{1+\delta}}\leq K,\quad[\omega]\cdot[\theta_{\widehat{Y}}]^{n-1}\leq A\ \text{ and }\ [\omega]^{n}\geq A^{-1},$
where $F_{\omega}:=\log(\frac{1}{V_{\omega}}\frac{\omega^{n}}{\theta_{\widehat{Y}}^{n}})$ for $V_{\omega}:=[\omega]^{n}$. For any $q\in(1,\frac{n}{n-1})$, there exists a constant $C=C(\gamma,A,K,\delta,q)>0$ such that for any $\omega\in\mathcal{K}(\widehat{Y},\theta_{\widehat{Y}},\gamma,A,K,\delta)$, we have
$(\int_{\widehat{Y}}|u|^{2q}\omega^{n})^{\frac{1}{q}}\leq C\int_{\widehat{Y}}(|du|_{\omega}^{2}+|u|^{2})\frac{\omega^{n}}{n!}$
@@
Building on Theorem 3.10, Proposition 3.9 can be directly concluded by applying the smooth approximation argument from [28, Section 3]. For the reader’s convenience, we include a brief explanation.
@@
For simplicity, we omit the subscript ‘k’ for now. Note that $\mu_{\alpha}:U_{\alpha}\rightarrow U_{\alpha}/G_{\alpha}$ is étale outside a proper analytic subspace $Z$ of codimension at least $2$. Recall from Proposition 2.19 that there exists a smooth Kähler form $\omega_{0}$ on the underlying space $Y$ such that $\omega_{Y_{\mathrm{orb}}}:=\{\mu_{\alpha}^{*}\omega_{0}+\sqrt{-1}\partial\bar{\partial}\varphi_{\alpha}\}$ for some $G_{\alpha}$-invariant smooth function $\varphi_{\alpha}$ on every chart $U_{\alpha}$. Then $\varphi_{\alpha}$ induces a bounded continuous function $\varphi$ on $Y$ that is smooth outside $Z$ and in particular, $\omega_{Y_{\mathrm{orb}}}=\{\mu_{\alpha}^{*}\omega_{Y}\}$ for $\omega_{Y}=\omega_{0}+\sqrt{-1}\partial\bar{\partial}\varphi$. Thus, $\omega_{\mathrm{orb},\epsilon}:=\{f_{\mathrm{orb}}^{*}\omega+\epsilon(\mu_{\alpha}^{*}\omega_{0}+\sqrt{-1}\partial\bar{\partial}\varphi_{\alpha})\}$ can also be viewed as a Kähler current $\omega_{\epsilon}=f^{*}\omega+\epsilon\omega_{0}+\epsilon\sqrt{-1}\partial\bar{\partial}\varphi$ on the underlying complex space $Y$ and $\omega_{\epsilon}$ is smooth outside $Z$. Let $g:\widehat{Y}\rightarrow Y$ be a resolution of singularities of $Y$ and $D$ be the exceptional divisor of $f\circ g$. We use the notation of Proposition 3.9 and Theorem 3.10.
As in [45, Section 7], the basic strategy is to consider a smooth approximation of $g^{*}\omega_{\epsilon}$ on $\widehat{Y}$. The situation differs slightly from that in [45, Section 7] because $\log(\frac{g^{*}\omega_{\epsilon}^{n}}{\theta_{\widehat{Y}}^{n}})$ does not, in general, have log type analytic singularities (see [45, Definition 7.2]). In fact, it is the sum of a function $F_{\log}$ with log type analytic singularities and a continuous function $G_{\epsilon}$ that is smooth outside $D$ (cf. [28, Lemma 3.9]). The main consideration in [28, Section 3] is addressing this issue. According to [28, Lemma 3.12, Lemma 3.13 and Lemma 3.15], there exists a sequence $\omega_{\epsilon,j}$ of Kähler metrics on $\widehat{Y}$ such that $\omega_{\epsilon,j}\in\mathcal{K}(\widehat{Y},\theta_{\widehat{Y}},\gamma,A,K,\delta)$ for some $A,K,\delta$ and $\gamma$ independent of $j$ and $\epsilon$, and for any integer $k\geq 0$, we have $\omega_{\epsilon,j}\rightarrow g^{*}\omega_{\epsilon}$ in $C_{loc}^{k}(\widehat{Y}\setminus D)$. Then by applying the argument on [45, Page 34] and Theorem 3.10, the uniform Sobolev inequality (3.16) for $\omega_{\mathrm{orb},\epsilon}$ can be proved because $\mu_{\alpha}$ is unramified outside $Z$. ∎
Remark 3.11.
It might be possible to develop an orbifold analogue of [45] to establish Proposition 3.9; see e.g. [97] for the orbifold version of the mean value inequality.
It is easy to see that there exists some constant $C_{k}>0$ independent of $\epsilon$ such that
@@
from the construction (see (3.5) and (3.6)). We have the following uniform Sobolev inequalities for the Hermitian metrics $\omega_{\mathrm{orb},\epsilon}$ by applying Proposition 3.9 to $q=\frac{2n}{2n-1}$.
@@
(3.18) $(\int_{Y_{\mathrm{orb}}}|u|^{\frac{4n}{2n-1}}\frac{\omega_{\mathrm{orb},\epsilon}^{n}}{n!})^{\frac{2n-1}{2n}}\leq C_{S}\int_{Y_{\mathrm{orb}}}(|du|^{2}_{\omega_{\mathrm{orb},\epsilon}}+|u|^{2})\frac{\omega_{\mathrm{orb},\epsilon}^{n}}{n!}$